### v1

# Evaluating The Explainability of State-of-the-Art Machine Learning-based IoT Network Intrusion Detection Systems

Abstract
Internet-of-Things (IoT) Network Intrusion Detection Systems (NIDSs) which use machine learning (ML) models achieve high detection performance and accuracy while avoiding dependence on fixed signatures extracted from attack artifacts. However, there is a noticeable hesitance among network security experts and practitioners when it comes to deploying ML-based NIDSs in real-world production environments due to their black-box nature, i.e., how and why the underlying models make their decisions. In this work, we analyze state-of-the-art ML-based IoT NIDS models using explainable AI (xAI) techniques (e.g., TRUSTEE, SHAP). Using the explanations generated for the models’ decisions, the most prominent features used by each NIDS model considered are presented. We compare the explanations generated across xAI methods for a given NIDS model as well as the explanations generated across the NIDS models for a given xAI method. Finally, we evaluate the vulnerability of each NIDS model to inductive bias (artifacts learnt from training data). The results show that: (1) some ML-based IoT NIDS models can be better explained than other models, (2) xAI explanations are in conflict for most of the IoT NIDS models considered in this work and (3) some IoT NIDS models are more vulnerable to inductive bias than other models.
Index Terms:
Network Intrusion Detection Systems, NIDS, Internet-of-Things, IoT, Machine Learning, explainable Artificial Intelligence, xAI
I Introduction
The Internet-of-Things (IoT) [1] constitutes a vast network of sensor-equipped devices, often characterized by low power and modest processing capabilities. These devices seamlessly exchange data among themselves or with central systems such as gateways or cloud servers using a mix of wired (such as Ethernet) and wireless technologies (such as RFID, Zigbee, WiFi, Bluetooth, and 3G/4G). The applications of IoT devices span a wide spectrum: from wearables like fitness trackers and smartwatches, to home automation systems managing lighting, thermostats, cameras, and door locks, all the way to industrial automation for tasks like manufacturing equipment management, process control, and ensuring plant safety. By 2032, experts estimate that a staggering 34.4 billion IoT devices will be deployed globally [2]. Unfortunately, cyber attackers have set their sights on IoT devices, leveraging malicious software (commonly known as malware) [3, 4]. Their motivations stem from several factors: first, the prevalence of legacy devices still connected to the Internet without adequate security updates; second, the tendency to prioritize other aspects over security during the development cycle; and third, the widespread use of weak login credentials. The sheer volume of deployed devices ranging from cameras to sensors, poses a formidable challenge when it comes to implementing cost-effective, real-time anomaly detection at such a massive scale.
Securing IoT networks against intrusion is paramount, and network intrusion detection systems (NIDS) play a pivotal role in achieving this. Traditionally, NIDSs have relied heavily on signature-based methods [5, 6]- essentially matching attack patterns to known signatures. However, these approaches demand significant manual effort from experts to create and maintain those signatures. Moreover, they fall short when it comes to identifying novel, previously unseen attacks. With the introduction of machine learning (ML)-based NIDSs which leverage ML models to enhance existing defenses, impressive performance and accuracy has been achieved [7, 8]. Unlike their signature-centric counterparts, ML-based NIDSs do not require experts to painstakingly prepare signatures in advance. In fact, unsupervised ML-based NIDSs can even raise alerts for hitherto unknown attacks. Thanks to these advantages and the rapid evolution of ML techniques, ML-based NIDSs have made significant strides in recent years.
Despite the promise of ML-based NIDSs, there is a noticeable hesitance among network security experts and practitioners when it comes to deploying them in real-world production environments [9]. The crux of the issue lies in the black-box nature of many of these proposed solutions. Essentially, their inner workings (how and why they make decisions) remain elusive, unlike simpler but less effective rule-based approaches that security professionals are accustomed to. Security analysts rely on ML-based NIDSs for critical decisions. However, without clear explanations for the decisions made by these underlying ML models, it becomes difficult for the analysts to trust them. Moreover, ML often operates as a black box within the NIDS. Consequently, hidden correlations stemming from unrelated artifacts can slip past undetected, leading to a flawed perception of the system’s capabilities. Explanation tools and techniques can unveil those hidden correlations, allowing practitioners to assess their impact on the NIDS’s overall effectiveness. As it stands, the current state-of-the-art NIDSs for IoT environments lack inherent explainability. Therefore, there is a pressing need to scrutinize them using these very explanation techniques.
In this work, we analyze state-of-the-art ML-based IoT NIDS models using explainable AI (xAI) techniques (e.g., TRUSTEE [10], SHAP [11]). Except Kitsune, other NIDS source codes do not provide trained ML models or training/test data in a format which can be directly input to the xAI methods’ APIs. Therefore, we make changes in the NIDS source codes to satisfy the APIs’ requirements. Using both global explanations and local explanations for the models’ decisions, the most prominent features used by each ML-based IoT NIDS model considered are presented. We compare the explanations generated across xAI methods for a given NIDS model (using our devised qualitative approach) as well as the explanations generated across the NIDS models for a given xAI method (using standard evaluation metrics). Finally, we evaluate the vulnerability of each NIDS model to inductive bias (artifacts learnt from training data). To the best of our knowledge, our work is the first one to generate explanations for state-of-the-art ML-based IoT NIDSs other than Kitsune using xAI techniques.
The main contributions of our work are as follows:
  • We analyze state-of-the-art ML-based IoT NIDS models using explainable AI techniques.
  • We present the most prominent features used by each IoT NIDS model considered to make decisions on input data samples using the explanations generated earlier.
  • We compare the explanations generated across the models for a given xAI method as well as the explanations generated across xAI methods for a given model.
  • We evaluate the vulnerability of each IoT NIDS model to inductive bias.
The rest of the paper is organized as follows. In Section II, we review published research on the need for explanation of ML systems targeted at cybersecurity, evaluation of explanation methods on ML-based security systems and explainable ML-based attack detection. In Section III, we present a background on the types of explanation methods for ML models, prominent black-box xAI methods, and an overview of the state-of-the-art ML-based IoT NIDSs to which we apply explanation techniques in this work. In Section IV, we explain the issues with applying xAI methods to the ML-based IoT NIDSs directly using their open-source implementations, our approach to addressing those issues and details of our investigation beyond direct application of xAI methods which covers the relation between explanations generated by different xAI methods and the detection of inductive bias in the IoT NIDSs under consideration in this work. Finally, in Section V, we evaluate the explanations generated for each of the IoT NIDSs using xAI methods, compare the explanations across the NIDS models as well across xAI methods, and discuss insights gleaned from the main results obtained thus far.
II Related Work
Several works have recommended employing explanation techniques for ML systems deployed for cybersecurity. In [9], the authors have recommended employing explanation techniques to delve deeper into the features of learning-based security systems. Not withstanding their limitations, these techniques can unveil spurious correlations and empower experts to evaluate their effect on the security system’s features. The authors in [12] have also emphasized the importance of explaining the outcomes produced by ML-based detectors used in security contexts. Beyond mere predictions, a cybersecurity threat detection model can offer valuable insights. By understanding the model’s explanations, security practitioners can enhance protection around monitored assets in subsequent executions. Explainability varies in relevance across different domains, but the authors contend that it is crucial for several cybersecurity tasks, enabling the effective application of countermeasures against security threats. The participants (security practitioners) of a survey conducted by the authors of [13] have expressed the view that security solution providers ought to prioritize methods that offer clear explanations to their clients.
As a result, we can see the beginning of efforts in the research community to apply xAI techniques to ML-based security systems. A few works [14, 15, 16, 17] have employed xAI methods such as LIME [18] and SHAP [11] to assess the quality of local/global explanations generated for conventional machine learning classifiers (AdaBoost, k-nearest neighbour, Multi-layer Perceptron, Random Forest, XGBoost, Light Gradient-Boosting Machine, Gradient Boosting Classifier) applied to IoT botnet datasets (N-BaIoT [19], MedBIoT [20], and BoT-IoT [21]) and network intrusion datasets (NSL-KDD dataset [22], CICIDS-2017 [23]) in terms of criteria such as faithfulness, stability, complexity, and sensitivity. In [24], the authors have delved into six explanation methods, evaluating their effectiveness across four security systems described in existing literature. These systems leverage deep learning techniques to detect Android malware, malicious PDF files and security vulnerabilities. The assessment criteria encompass both general properties of deep learning and domain-specific aspects relevant to security.
The authors in [25] have introduced a framework rooted in SHAP to provide explanations for IDSs. This framework blends both local and global explanations, enhancing the overall interpretability of IDSs. Local explanations shed light on why the ML model employed by a specific IDS reaches particular decisions for individual inputs. Meanwhile, global explanations highlight crucial features extracted from the ML model and elucidate the relationships between feature values and specific attack types. [26] has presented a deep learning-based IDS for IIoT networks consisting of an ensemble of three CNN models and an extreme-learning machine model. The authors have used SHAP and LIME techniques to explain the ensemble detector’s decision-making process.
A few works have also proposed ML-based attack detection systems with built-in explainability. In [27], the authors have implemented an xAI solution to provide accurate detection and classification of the DNS-over-HTTPS attacks. It is based on a balanced stacked random forest classifier. The authors have highlighted the underlying feature contributions to provide transparent and explainable results from the model. In [28], the authors have proposed an NIDS for IT networks combining ensemble learning and stacking with a meta-learner (CNN) that works on graphical representation of traffic flows and provides the required explainability level for the decisions made. They also provide visual representations of network anomalies that allows security analysts to interpret and gain insights into the detected network anomalies.
Our work differs from the previous literature in the following aspects:
  • Existing works that evaluate xAI methods on ML-based IoT NIDSs do not focus on state-of-the-art IoT NIDSs and do not include the latest xAI tools for ML-based network security such as TRUSTEE ([14, 15, 16, 17]).
  • Existing explainable-by-design AI-based attack/intrusion detection systems are targeted at specific attacks (DNS-over-HTTPS attacks in [27]) and non-IoT networks (IT networks in [28]).
III Background
In this paper, since we delve into explaining the decisions of state-of-the-art ML-based IoT NIDSs, it is essential to briefly explore the types of explanation methods present in academic literature, the IoT NIDSs that we consider for explanation in this work and the various aspects of our investigation with xAI techniques on those IoT NIDSs.
III-A Explanation Methods- Types
All explanation methods for ML models can be broadly categorized into black-box and white-box explanations.
Black-box Explanations These methods function within a black-box context, assuming zero knowledge about the inner workings of the underlying ML model and its associated parameters. These methods prove invaluable when direct access to the ML model is not available, for instance, during remote audits of a learning-based application. These black-box methods approximate the underlying ML model’s prediction function $f_{N}$, allowing them to estimate how various dimensions of the input vector $x$ affect a given prediction. While they hold promise for explaining deep learning models, their effectiveness can be impeded by the black-box environment, sometimes leaving out crucial insights embedded in the ML model’s parameters and architecture.
White-box Explanations These methods work under the premise that we have full knowledge of all the parameters of the ML model being tested. Armed with this information, these methods bypass approximations and can directly compute explanations for the prediction function, $f_{N}$. The same ML model is often used to generate both predictions and explanations, assuming that the model is available on hand. This assumption holds true for stand-alone automated systems such as those that analyse binaries, detect malware or discover software vulnerabilties. However, it is worth noting that some white-box methods are customized for specific neural network architectures deployed in the fields of computer vision/speech processing/natural language processing and may not perform well for other architectures.
III-B Prominent Black-box XAI Tools
ML-based IoT NIDS products are expected to be used by end users/customers who do not have access to the ML model and its parameters underlying the NIDS. This is to maintain confidentiality of the NIDS implementation by the cybersecurity services vendor. Therefore, we assume that the ML-based IoT NIDSs being explained in this work are available to us as a black-box system only and we use xAI techniques designed to explain such black-box ML systems. In what follows, we present a brief overview of each black-box xAI tool that we have tested with as part of this work.
Overview of TRUSTEE: TRUSTEE [10] is a state-of-the-art interpretability tool for ML-based security solutions which generates interpretable decision trees (DT) for a given ML model. It can be used to find out inductive biases in black-box models (e.g., shortcut learning, spurious correlations, sensitiveness to out-of-distribution samples). While generating DT, TRUSTEE focuses on the following desirable properties: model-agnosticism (should be applicable to any given black-box model), model fidelity (how accurate the DT is with respect to the original ML model), model stability (the decision rules should correctly describe how the given black-box model makes a significant number of its decisions and should be largely insensitive to the particular data samples that TRUSTEE used in the process of selecting its final DT explanation) and model complexity (decision rules should be readily recognizable by domain experts). The important aspects of generated DT explanation are summarized in a trust report which helps end users determine whether the given black-box model suffers from the problem of underspecification and can be trusted.
Overview of LIME: Ribeiro et al. [18] introduced LIME, one of the first black-box methods for explaining neural networks that is further extended by SHAP [11]. Both works are motivated by the need to allow users to be able to trust ML models as well as their predictions. LIME is designed to explain the predictions of any classifier or regressor in a faithful way, by approximating it locally with an interpretable model. The main characteristics of LIME are:
  • Interpretable, i.e., it provides qualitative understanding of the relation between the input variables and the response.
  • Locally faithful, i.e., it corresponds to how the model behaves in the vicinity of the instance being predicted.
  • Model-agnostic, i.e., it treats the original ML model as a black box.
LIME aims at approximating the decision function, $f$ of the model being explained by solving the following optimization problem:
$\underset{g\in\mathcal{G}}{argmin}\ \mathcal{L}(f,g,\pi_{x}),$ (1)
where $\mathcal{L}(f,g,\pi_{x})$ is the fidelity function which is a measure of how unfaithful the interpretable model $g$ is in approximating $f$ in the locality defined by $\pi_{x}$, while $\pi_{x}(z)$ is the proximity measure between an instance $z$ to $x$, so as to define the locality around $x$ and $x$ is the input to $f$.
Overview of SHAP: SHAP [11] (SHapley Additive exPlanations) subsumes additive feature attribution methods such as LIME and uses the following forms for $\mathcal{L}(f,g,\pi_{x})$ and $\pi_{x}$ under the SHAP kernel method, which is shown to recover Shapley Values [29] when solving the regression:
$\mathcal{L}(f,g,\pi_{x^{\prime}})=\sum_{z^{\prime}\in Z}[h_{x}^{-1}(z^{\prime})-g(z^{\prime})]^{2}\pi_{x^{\prime}}(z^{\prime}),$
$\pi_{x^{\prime}}(z^{\prime})=\frac{M-1}{(M\text{choose}|z^{\prime}|)|z^{\prime}|(M-|z^{\prime}|)}$
where $x^{\prime}$ is the simplified input which maps to the original input $x$ through a mapping function $x=h_{x}(x^{\prime})$, $M$ is the number of simplified input features, $z^{\prime}\in{\{0,1\}}^{M}$, and $|z^{\prime}|$ is the number of non-zero elements in $z^{\prime}$.
Shapley values are a concept from game theory where the features act as players under the objective of finding a fair contribution of the features to the payout, in this case the prediction of the model. Different from TRUSTEE, Shapley values are used for local interpretation. In other words, it is used to interpret how a prediction result is reached for a specific data sample or subset of samples. Specifically, Shapley values can tell how each feature contributes to the predicted results. A positive Shapley value means that the feature’s value pushes the classification result toward being malicious, and a negative Shapley value does the contrary.
Overview of LEMNA: LEMNA [30] is a black-box method specifically designed for security applications. Given an input data instance $x$ and a classifier such as an RNN, this method aims to identify a small set of features that have key contributions to the classification of $x$. This is done by generating a local approximation of the target classifier’s decision boundary near $x$. It uses a mixture regression model (to approximate locally non-linear decision boundaries) enhanced by fused lasso (to handle correlated features) for approximation. The mixture regression model is a weighted sum of $K$ linear regression models:
$f(x)=\sum_{j=1}^{K}\pi_{j}(\beta_{j}\cdot x+\epsilon_{j}).$ (2)
The parameter $K$ specifies the number of models, the random variables $\epsilon=(\epsilon_{1},\dots,\epsilon_{K})$ originate from a normal distribution $\epsilon_{i}\sim\mathcal{N}(0,\sigma)$ and $\pi=(\pi_{1},\dots,\pi_{K})$ holds the weights for each model. The variables $\beta_{1},\dots,\beta_{K}$ are the regression coefficients and can be interpreted as $K$ linear approximations of the decision boundary near $f(x)$.
xAI Evaluation Metrics: Several metrics have been proposed in literature to quantify the effectiveness of xAI techniques. The authors in [31] have categorized existing XAI evaluation metrics into subjective and objective metrics. Subjective metrics rely on human feedback, e.g., comprehensibility, satisfaction, and persuasiveness. This feedback can come from randomly selected users or domain experts. Objective metrics are grounded in formal definitions, e.g., explanation completeness which quantifies the extent to which a prediction’s contributing features are captured, and perturbation metrics which assess sensitivity to input changes by comparing outputs and considering model parameter variations across different instances of the same model.
III-C State-of-the-art ML-based IoT NIDSs
We now present a brief overview of each state-of-the-art ML-based IoT NIDS that we investigate using xAI techniques in this work.
Kitsune [32] proposes an online unsupervised anomaly detection system based on an ensemble of autoencoders which is lightweight in terms of memory footprint and meant to be deployed on network gateways and routers to detect attacks on the local network. It uses a damped incremental statistics method to extract network traffic features which has the advantage of extracting features from dynamic network traffic at high speed.
HorusEye [33] is a two-stage anomaly detection framework for IoT. In the first stage, preliminary burst-level anomaly detection is implemented on the data plane. The suspicious traffic is then reported to the control plane for further investigation. To reduce the false positive rate, the control plane carries out the second stage, where more thorough anomaly detection is performed over the reported suspicious traffic using flow-level features and an autoencoder.
ENIDrift [34] addresses real-world issues in ML-based NIDS deployment such as concept drift (change in statistical properties of target variable over time), class-imbalanced training data and well-crafted ML attacks (training data contamination, adversarial attacks). It includes three main components: iP2V- an incremental feature extraction method based on Word2Vec, a sub-classifier generation module and an update module.
HELAD [35] uses ensemble learning consisting of an unsupervised autoencoder as a base learner which learns the profile for normal network traffic and a supervised LSTM (Long-term short-term memory) which is trained using the autoencoder’s output to detect continuous attacks. The model is re-trained using a new concept of time slice re-training. It also uses dynamic thresholds and integrated learning parameters to keep the anomaly detection performance from degradation.
IoTA [36] is an intrusion detection system for IoT devices which uses fully packet-level models to profile IoT device traffic patterns. These models are based on automata constructed from raw device traffic traces whose states represent packets with similar context, representative features are packet length and direction and when grouped together based on connectedness, represent device traffic flows. At runtime, ongoing IoT device traffic is matched with the extracted normal traffic profiles towards intrusion detection. However, the authors have not released the source code for IoTA implementation publicly and therefore, we have not included it in our analysis.
IV xAI Analysis of ML-based IoT NIDSs
In this work, we analyze four state-of-the-art ML-based IoT NIDSs (Kitsune, HorusEeye, ENIDrift, HELAD) using xAI techniques. However, xAI techniques such as TRUSTEE and kernel SHAP can not be applied directly to the IoT NIDSs under consideration. This is because xAI methods have certain requirements at the API level which are not fulfilled by the current implementations of the NIDSs. We had to modify the NIDS implementations to make them compatible with xAI methods. Still, there were a few xAI methods whose requirements were found to be incompatible with the NIDS implementations unless we made major changes to the source code. Additionally, we did not limit ourselves to a direct application of xAI methods to the NIDSs but we also investigated: (1) the relation between the explanations generated for a given NIDS by different xAI methods, and (2) the presence of inductive biases in the NIDS implementations.
IV-A xAI-compatible IoT NIDS Implementation
The TRUSTEE trust report generating API has two requirements: (1) a trained ML model with a predict() function taking training/test dataset as the only input and outputting predicted labels, and (2) training/test datasets and corresponding predicted labels should be in numpy array/pandas dataframe format. The kernel SHAP API shares the requirements of TRUSTEE API. Except Kitsune, other NIDS source codes do not provide trained ML model or training/test data in a format which can be directly input to TRUSTEE API or kernel SHAP API.
For example, in the case of ENIDrift source code, the input dataset and predicted labels were not in numpy array/pandas dataframe format while the predict() function was not designed for an input dataset in the form of a 2-D matrix with rows as samples and columns as feature values. Further, ENIDrift does not extract explicit features from incoming packets but instead embeds packets to vectors which may not be meaningful for model explainability analysis. In the case of HorusEye source code, the main script was in the form of a command with argument parsing, all library functions were defined in the same main script, there was no predict() function and input dataset and predicted labels were not in numpy array/pandas dataframe format. In the case of HELAD source code, there was no predict() function implemented and the NIDS model returned anomaly scores instead of class labels. Further, the way anomaly scores were calculated was not compatible with TRUSTEE API implementation.
While running LIME analysis of above NIDSs, we received an error that LIME does not support ML models which do not have $predict\_proba()$ function implemented. Since all the state-of-the-art ML-based IoT NIDSs considered in this paper do not implement $predict\_proba()$ function, we are forced to leave out LIME analysis of the NIDSs from this paper. Here, $predict\_proba()$ function returns the predicted probabilities of the input features belonging to each category. This method, instead of returning a discrete class, returns the probabilities associated with each class. This is useful when we not only want to know the predicted category of the input features, but also the model’s confidence in its prediction.
While running LEMNA analysis of above NIDSs, we found that the current implementation (last changed 5 years ago) only supports basic MLP, CNN and RNN models. Custom ML models such as the IoT NIDS models which are under consideration in this work are not supported. Further, LEMNA does not output feature-based model explanations for processing by cybersecurity experts such as those offered by TRUSTEE and SHAP.
IV-B Investigation Beyond Direct xAI Analysis
Relation between TRUSTEE and SHAP Interpretations: To establish the relationship between TRUSTEE global explanations and the SHAP local explanations of the black-box model underlying a given IoT NIDS, we can follow two approaches:
  • Naïve approach- We can directly compare the kernel SHAP features with highest SHAP values against the top-k features identified by TRUSTEE report. However, this comparison would not be meaningful as the SHAP value for each feature is calculated for a subset of data samples and reflects the importance of that feature in the decision made by the black-box model for a majority of the data samples in that subset. On the other hand, TRUSTEE acts on the complete input dataset and identifies the top features contributing the most to forming the decision boundary separating all benign and malicious data samples which are part of the input dataset.
  • Systematic approach- Since xAI methods such as SHAP are more suited towards analyzing a subset of data samples [10], they can be applied to the subset of samples corresponding to a path in a TRUSTEE DT starting at the root node and ending at one of the leaf nodes. We can compare the explanation thus generated with the one provided by TRUSTEE.
Under the systematic approach, for a given subset of data samples $D$, if $T$ is the set of features used in the selected root node-leaf node DT path decisions and $S$ is the set of features identified by kernel SHAP that contribute most to the classification decision for majority of the samples in $D$, then $T\subset S$ should hold true for the TRUSTEE DT explanation to agree with the SHAP explanation of the black-box model at the subset level. To form the set of features $S$, the mean absolute value of kernel SHAP values for multiple data samples are calculated for every feature, and subsequently the features are ranked according to their aggregate SHAP values to find the top features.
Inductive Bias: A number of ML models face the issue of underspecification. This term refers to the challenge of determining whether a model’s high performance is due to its genuine ability to capture essential patterns in the data or merely the result of certain inductive biases it has learned. These biases often appear as spurious correlations, a lack of ability to adapt to out-of-distribution (o.o.d.) data or shortcut learning strategies. The presence of such biases in trained models undermines their reliability and expected performance in real-world applications. Therefore, identifying these inductive biases is crucial for establishing trust in ML models. To detect the presence of an inductive bias in the ML-based IoT NIDSs under consideration in this work, such as vulnerability to o.o.d. samples, we use tampered malicious traffic traces [10] which alter the attack segments of the original traffic traces by distributing the malicious packets in such a way that the packet rate remains within a specified limit.
V Experimental Results
In this section, we present explanations of the four state-of-the-art ML-based IoT NIDSs (Kitsune, HorusEeye, ENIDrift, HELAD) generated using xAI methods such as TRUSTEE and kernel SHAP. For each pair of NIDS and xAI method, we wrote Python scripts for generating the explanation using the xAI method’s API and the modified source code of the NIDS (to make it compatible with the xAI method’s API implementation). Using the generated explanations, we present the most prominent features used by each NIDS model to make decisions on input data samples. We also investigate the relation between the explanations generated for a given NIDS by TRUSTEE and kernel SHAP by using the systematic approach for comparing TRUSTEE and SHAP explanations at the DT subset-level as explained in sub-section IV-B. Subsequently, we investigate the presence of inductive biases (e.g., vulnerability to o.o.d. data samples) in the NIDS implementations using tampered malicious traffic traces as explained in sub-section IV-B. Finally, we compare the explanations generated across the four NIDSs by TRUSTEE using the fidelity metric and by kernel SHAP using general xAI evaluation metrics.
Datasets: We use the Kitsune Mirai dataset [32] and the CICIDS-2017 dataset [23] for training/testing the ML-based IoT NIDSs under consideration. The Mirai trace consists of $\sim$ 760k packets of synthetically generated attack in a network with nine IoT devices, first $\sim$120k packets consist of benign traffic, remaining $\sim$640k packets have anomalous traffic. The CICIDS-2017 dataset [23] consists of five days of packet capture (July 3 to July 7, 2017) from a small-scale enterprise-like network, with the first day consisting of benign background traffic and the other days consisting of attack traffic (brute force SSH, DoS, Heartbleed, web attack, infiltration, etc.). It consists of significant drift over the data collection period in terms of different attack tools used. Below, we present the results using Mirai dataset only. Since the results obtained with CICIDS-2017 dataset are similar to those obtained with the Mirai dataset, we include them in the Appendix Section A.
Test Environment: To test Kitsune, HorusEye and ENIDrift NIDSs with xAI methods, we used a VMWare ESXi server VM with Intel Xeon Silver 4216 CPU @2.10GHz, 64-bit architecture, 8 cores, 16GB RAM and running Ubuntu 18.04/Ubuntu 20.04 OS. For testing HELAD NIDS, we used an Nvidia RTX 3090 GPU cluster with each GPU @1.4GHz, $\sim$25GB of memory since it requires training deep belief networks and LSTMs. Each GPU had 328 tensor cores and was running Ubuntu 22.04 OS.
V-A Case Study- Kitsune $\times$ Mirai
TRUSTEE Analysis: The results of TRUSTEE analysis for Kitsune are shown in Table I. The sample size refers to the fraction of the training dataset to use to train the student decision tree model. Fidelity is defined as the R-squared value between Kitsune’s predictions and those obtained by the DT explanation. Using TRUSTEE with $30\%$ of the original Mirai dataset samples and no pruning results in the DT explanation which achieves 0.786 fidelity compared to Kitsune. Using TRUSTEE with $30\%$ of the dataset samples and top-k pruning method (setting $k=3$) results in a DT explanation that is shown in Fig. 1 and achieves 0.697 fidelity compared to Kitsune.
Using TRUSTEE with $30\%$ of the dataset samples and top-k pruning method (setting $k=10$) results in a DT explanation which achieves 0.721 fidelity compared to Kitsune. The reduction in fidelity with pruning branches of a DT is expected since the obtained smaller trees run the risk of missing important decision branches as they prevent the consideration of any further decision branches once the stopping criterion is reached.
The top three prominent features that Kitsune uses to determine an anomaly are:
  • $HpHp\_0.01\_pcc\_0\_1$, $HpHp\_0.1\_pcc\_0\_1$- Correlation coefficient between two packet size streams aggregated by traffic sent between a set of source and destination IP addresses with time windows 0.01 (1 minute) and 0.1 (10 seconds)
  • $HpHp\_0.1\_weight\_0$- Weights (packet count) aggregated by traffic sent between a set of source and destination IP addresses and associated with the time window 0.1 (1.5 seconds).
Thus, Kitsune’s DT relies mainly on the volume of packets and sizes of packets exchanged per time frame to determine if an attack is underway.
Sample size Top-k pruning used? DT size, depth, leaves Fidelity
30% No 133377, 70, 66689 0.786
30% Yes (k=3) 19, 9, 10 0.697
30% Yes (k=10) 37, 13, 19 0.721
TABLE I: TRUSTEE analysis results for Kitsune Mirai model
Figure 1: Decision tree for Kitsune Mirai model with Top-k pruning (k=3). Only the top 3 layers are shown.
SHAP Analysis: Fig. 2 illustrates a local explanation for a benign packet data point using a SHAP’s force plot for Kitsune, displaying the contribution of each feature to the prediction. The plot shows the base value, and the features containing a positive influence on the prediction are in red, and the features showing a negative influence on the predictions are in blue. The base value in the plots is the average of all prediction values. Each strip in the plot illustrates how the features influence the predicted value, either drawing it closer or pushing it farther away from the base value. Red strip features push the value to higher values, whereas blue strip features push the value to lower values. The contribution of features holding broader strips is more.
The base value is $0.0646$. Features such as $HpHp\_0.1\_weight\_0$ and $MI\_dir\_0.1\_weight$ have a negative impact on the prediction value. $HpHp\_0.1\_weight\_0$ is the most crucial feature, as the contribution has a broader range. The total negative contribution is greater than the positive contribution, and the final predicted value is lesser than the base value. As a result, the class is predicted as benign. The feature $HpHp\_0.1\_weight\_0$ corresponds to weights aggregated by traffic sent between a set of source and destination IP addresses using a specific protocol (ARP/TCP/UDP) with time window 0.1 (10 seconds) while the feature $MI\_dir\_0.1\_weight$ corresponds to weights aggregated by source MAC and IP addresses, with time window 0.1 (10 seconds).
Figure 2: SHAP force plot for a single data point from Kitsune Mirai model
Using kernel SHAP analysis, Fig. 3 shows a summary of how the top features in Mirai dataset impact the Kitsune NIDS model’s output through a beeswarm plot for randomly selected 1000 data samples corresponding to the benign root node-leaf node path in the TRUSTEE DT (Fig. 1) which takes $625.70$ seconds on an average. Running SHAP analysis for the whole dataset will take substantial amount of time and computing resources. Each instance of the given explanation is represented by a single dot on each feature row. The $x$ position of the dot is determined by the SHAP value of that feature, and dots pile up along each feature row to show density. Color is used to display the original value of a feature. The features are ordered using the mean absolute value of the SHAP values for each feature. Features such as $MI\_dir\_0.1\_weight$, $HH\_0.01\_weight\_0$ and $HH\_0.1\_weight\_0$ have the most impact on the prediction value, in that order. $HH\_0.01\_weight\_0$ and $HH\_0.1\_weight\_0$ correspond to weights aggregated by traffic sent between a set of source and destination IP addresses with time windows 0.01 (1 minute) and 0.1 (10 seconds) respectively.
TRUSTEE vs SHAP Analysis: Using the systematic approach for comparing TRUSTEE and SHAP explanations as explained in sub-section IV-B, since the set of features used in the benign root node-leaf node path in the TRUSTEE DT is not a subset of the set of top contributing features computed by kernel SHAP summary analysis ($HH\_0.1\_weight\_0$ is the only common feature between the two sets), TRUSTEE AND SHAP explanations do not agree with each other at the subset level for Kitsune NIDS.
Figure 3: SHAP beeswarm plot for Kitsune Mirai model
Detection of Inductive Bias: It was shown in [10] that Kitsune NIDS is vulnerable to o.o.d. samples when used with the Mirai dataset introduced in [32]. The authors evaluated the following threshold ranges (in packets per second) for ARP requests originating from devices infected with Mirai: (1) 10-50, (2) 30-70 and (3) 50-90. The original trace as well as the tampered traces are shown in Fig. 4. It can be observed that tampering had a significant effect on the difference between benign and attack traffic in terms of number of packets per second. Kitsune was executed on each trace with an identical number of training samples. The results showed that the RMSE values generated by Kitsune NIDS decreased significantly with a drop in the attack traffic volume, reaching the values corresponding to benign traffic.
Figure 4: Number of packets per second for original Kitsune Mirai trace and tampered traces. Blue portions of the plot correspond to benign traffic while red portions correspond to malicious traffic inclusive of benign and attack traffic.
V-B Case Study- HorusEye $\times$ Mirai
TRUSTEE Analysis: The results of TRUSTEE analysis for HorusEye are shown in Table II. Using TRUSTEE with $30\%$ of the Mirai dataset samples with no pruning results in a DT explanation which achieves negative fidelity ($\sim-1.365$) compared to original HorusEye model. Using TRUSTEE with $30\%$ of the dataset samples and top-k Pruning method (setting $k=10$) results in a DT explanation which achieves negative fidelity ($\sim-1.038e-06$) compared to original HorusEye model. Negative fidelity values mean that the DT explanation fits HorusEye predictions worse than the mean of the predictions.
Nevertheless, the decision tree for HorusEye Mirai model with top-k pruning (k=10) is shown in Fig. 5. The top three prominent features that HorusEye’s DT uses to determine an anomaly are:
  1. 1.
    $HH\_3\_pcc\_0\_1$- Correlation coefficient between two packet size streams aggregated by traffic sent between a set of source and destination IP addresses with time window 3 (500 milli-seconds).
  2. 2.
    $HH\_1\_std\_0$- Standard deviation of packet sizes aggregated by the traffic sent between a set of source and destination IP addresses with time window 1 (1.5 seconds).
  3. 3.
    $MI\_dir\_0.01\_mean$- Mean of packet sizes aggregated by source MAC and IP addresses, with time window 0.01 (1 minute).
Thus, HorusEye’s DT relies mainly on the sizes of packets exchanged between a pair of IP addresses per time frame to determine if an attack is underway.
Sample size Top-k pruning used? DT size, depth, leaves Fidelity
30% No 45, 11, 23 -1.365
30% Yes (k=10) 1, 0, 1 -1.038e-06
TABLE II: TRUSTEE analysis results for HorusEye Mirai model
Figure 5: Decision tree for HorusEye Mirai model with Top-k pruning (k=10). Only the top 3 layers are shown.
SHAP Analysis: Using kernel SHAP analysis, Fig. 6 shows a summary of how the top features in Mirai dataset impact the HorusEye NIDS model’s output using a beeswarm plot for randomly selected 1000 data samples corresponding to the benign root node-leaf node path in the TRUSTEE DT (Fig. 5). Features such as $HH\_0.1\_pcc\_0\_1$, $HH\_1\_mean\_0$ and $HH\_0.01\_std\_0$ have the most impact on the prediction value, in that order. $HH\_0.1\_pcc\_0\_1$ corresponds to the correlation coefficient between two packet size streams aggregated by traffic sent between a set of source and destination IP addresses with time window 0.1 (10 seconds), $HH\_1\_mean\_0$ corresponds to the mean of packet sizes aggregated by traffic sent between a set of source and destination IP addresses with time window 1 (1.5 seconds) and $HH\_0.01\_std\_0$ corresponds to the standard deviation of packet sizes aggregated by the traffic sent between a set of source and destination IP addresses with time window 0.01 (1 minute).
Figure 6: SHAP beeswarm plot for HorusEye Mirai model
TRUSTEE vs SHAP Analysis: Using the systematic approach for comparing TRUSTEE and SHAP explanations as explained in sub-section IV-B, the set of features used in the benign root node-leaf node path in the TRUSTEE DT is a subset of the set of top contributing features computed by kernel SHAP summary analysis. Therefore, TRUSTEE AND SHAP explanations agree with each other at the subset level for HorusEye NIDS.
Detection of Inductive Bias: Following a similar approach as the Kitsune case, three separate tampered traces with different attack traffic volumes were obtained from the Mirai dataset and HorusEye was executed on each trace. In Fig. 7(a), we show the results for first 200,000 packets of each traces. In Fig. 7(b), we also compare the expected RMSE (produced by HorusEye in the original trace) and the predicted RMSE for each tampered trace. The results show that the RMSE values generated by HorusEye decrease with a drop in the attack traffic volume. However, the difference between the RMSE values produced by the original and tampered traces is not as pronounced as in the case of Kitsune NIDS. This experiment demonstrates that the Mirai use case from HorusEye is vulnerable to o.o.d. samples, though to a lesser extent than Kitsune. It should be noted here that HorusEye [33] authors had recognized this issue (without an empirical analysis though) by stating that the frequency of Mirai attacks is too high in the original dataset to reflect real botnet infection. Therefore, the authors had randomly down-sampled the Mirai attack traffic in Kitsune dataset to consist of 40,000 packets during HorusEye’s performance evaluation.
(a) Predicted RMSE vs packet count
(b) Predicted vs expected RMSE (produced by HorusEye in the original trace)
Figure 7: HorusEye RMSE values for the first 200k packets from original and tampered traces.
V-C Case Study- ENIDrift $\times$ Mirai
TRUSTEE Analysis: The results of TRUSTEE analysis for ENIDrift are shown in Table III. Using TRUSTEE with $30\%$ of the Mirai dataset samples and no pruning results in the DT explanation which achieves 0.715 fidelity compared to ENIDrift. Using TRUSTEE with $30\%$ of the dataset samples and top-k Pruning method (setting $k=10$) results in a DT explanation (shown in Fig. 8) which achieves negative fidelity ($\sim-1.536e-09$) compared to ENIDrift. This is expected since pruning the branches of a DT reduces its fidelity. The most prominent features ENIDrift uses to determine anomaly are $f44$ and $f9$. ENIDrift does not extract explicit features from incoming packets. Instead, it uses ip2V incremental embedding technique based to embed packets to vectors. The embedding vector length in ENIDrift implementation is 200.
Sample size Top-k pruning used? DT size, depth, leaves Fidelity
30% No 5, 2, 3 0.715
30% Yes (k=10) 1, 0, 1 -1.536e-09
TABLE III: TRUSTEE analysis results for ENIDrift Mirai model
Figure 8: Decision tree for ENIDrift Mirai model with Top-k pruning (k=10).
SHAP Analysis: Fig. 9 shows a summary of how the top features in Mirai dataset impact the ENIDrift NIDS model’s output using a beeswarm plot for randomly selected 300 data samples. Features (embedding vector components) such as $f115$, $f89$, $f47$ have the most impact on the prediction value, in that order.
Figure 9: SHAP beeswarm plot for ENIDrift Mirai model
TRUSTEE vs SHAP Analysis: Using the systematic approach for comparing TRUSTEE and SHAP explanations as explained in sub-section IV-B, the set of features used in the benign root node-leaf node path in the TRUSTEE DT is not a subset of the set of top contributing features computed by kernel SHAP summary analysis (there are no common features between the two sets). Therefore, TRUSTEE AND SHAP explanations do not agree with each other at the subset level for ENIDrift NIDS.
Detection of Inductive Bias: Following a similar approach as the Kitsune case, three separate tampered traces with different attack traffic volumes were obtained from the Mirai dataset and ENIDrift was executed on each trace. In Fig. 10(a), we show the results for first 120,000 packets of each traces. In Fig. 10(b), we also compare the expected (weighted) ensemble scores (produced by ENIDrift in the original trace) and the predicted (weighted) ensemble scores for each tampered trace. The results show that the ensemble scores produced by ENIDrift do not have a discernable pattern of dependence on the volume of the attack traffic encountered. Since the distance of the dots for tampered traces from the diagonal line in Fig. 12(b) does not follow any pattern, the effect that tampering with the Mirai attack traces had on the predicted ensemble scores and hence, whether the Mirai use case from ENIDrift is vulnerable to o.o.d. samples or not, remains inconclusive.
(a) Predicted ensemble score vs packet count
(b) Predicted vs expected ensemble score (produced by ENIDrift in the original trace)
Figure 10: ENIDrift ensemble scores for the first 120k packets from original and tampered traces.
V-D Case Study- HELAD $\times$ Mirai
TRUSTEE Analysis: The results of TRUSTEE analysis for HELAD are shown in Table IV. Using TRUSTEE with $30\%$ of the Mirai dataset samples and no pruning results in the DT explanation which achieves 0.0 fidelity compared to HELAD. Using TRUSTEE with $30\%$ of the dataset samples and top-k Pruning method (setting $k=10$) results in a DT explanation (shown in Fig. 11) which achieves 0.0 fidelity compared to HELAD. A fidelity value of 0.0 means that the DT explanation fits HELAD predictions only as good as the mean of the predictions.
The top three prominent features HELAD’s DT uses to determine an anomaly are:
  1. 1.
    $MI\_dir\_3.0\_weight$- Weights aggregated by source MAC and IP addresses, with time window 3.0 (500 milli-seconds).
  2. 2.
    $HH\_1.0\_magnitude\_0\_1$- Absolute magnitude of two packet size streams aggregated by the traffic sent between a set of source and destination IP addresses with time window 1.0 (1.5 seconds).
  3. 3.
    $HH\_0.01\_weight\_0$- Weights aggregated by traffic sent between a set of source and destination IP addresses with time window 0.01 (1 minute).
Thus, HELAD’s DT relies mainly on the volume of packets and the sizes of packets exchanged between a pair of IP addresses per time frame to determine if an attack is underway.
Sample size Top-k pruning used? DT size, depth, leaves Fidelity
30% No 26289, 100, 13145 0.0
30% Yes (k=10) 435, 94, 218 0.0
TABLE IV: TRUSTEE analysis results for HELAD Mirai model
Figure 11: Decision tree for HELAD Mirai model with Top-k pruning (k=10). Only the top 3 layers are shown.
SHAP Analysis: Running a kernel SHAP analysis of HELAD NIDS with the Mirai dataset resulted in issues. First, we ran the analysis with 200,000 packets but it threw a numpy memory error since it was unable to allocate a huge amount of memory (335 GiB) for storing a ($2248\times 20000000$) array with 64-bit float values. Next, we ran the analysis with 20,000 packets but the LSTM was taking too long to train. Hence, we decided to not proceed with the kernel SHAP analysis.
Detection of Inductive Bias: Following a similar approach as the Kitsune case, three separate tampered traces with different attack traffic volumes were obtained from the Mirai dataset and HELAD was executed on each trace. In Fig. 12(a), we show the results for first 200,000 packets of each traces. In Fig. 12(b), we also compare the expected anomaly scores (produced by HELAD in the original trace) and the predicted anomaly scores for each tampered trace. The results show that the anomaly scores produced by HELAD for the original and tampered traces overlap with each other and therefore, the scores do not have a significant dependence on the volume of the attack traffic encountered. Since the dots for tampered traces are close to the diagonal line in Fig. 12(b), tampering with the Mirai attack traces had limited outcome on the predicted anomaly scores. This experiment demonstrates that the Mirai use case from HELAD is not vulnerable to o.o.d. samples to a significant extent.
(a) Predicted anomaly score vs packet count
(b) Predicted vs expected anomaly score (produced by HELAD in the original trace)
Figure 12: HELAD anomaly scores for the first 200k packets from original and tampered traces.
V-E Comparative Analysis
In this sub-section, we use xAI evaluation metrics to quantify the effectiveness of the explanations generated for the ML-based IoT NIDSs under consideration by TRUSTEE and kernel SHAP. We use those metrics to compare the explanations across NIDSs. Currently, the existing implementations of xAI evaluation metrics support kernel SHAP only and none of the implementations support TRUSTEE. TRUSTEE’s implementation comes with its own metric called fidelity which quantifies how accurate the TRUSTEE DT is with respect to the original black-box ML model.
TRUSTEE Fidelity: Using the Mirai dataset, if we compare the fidelity values obtained for the IoT NIDSs considered, without any pruning, the highest fidelity values are obtained for Kitsune (0.786) out of the four NIDSs. With top-k pruning ($k=10$), the highest fidelity values are obtained for Kitsune (0.721) out of the four NIDSs. Since ENIDrift, HorusEye and HELAD NIDSs yield negative and zero fidelity values, TRUSTEE DT explanations do not serve much purpose. Since top-k pruning is required to reduce the complexity of DT explanations and make them practical for cyber security personnel, TRUSTEE is most useful for explaining the decisions of Kitsune NIDS.
Using the CICIDS-2017 dataset, if we compare the fidelity values obtained for the IoT NIDSs considered (refer Appendix Section A), without any pruning, the highest fidelity values are obtained for ENIDrift (0.929) out of the four NIDSs. With top-k pruning ($k=10$), the highest fidelity values are obtained for Kitsune (0.919) out of the four NIDSs. Since HorusEye and HELAD NIDSs yield negative and zero fidelity values, TRUSTEE DT explanations do not serve much purpose. With top-k pruning, TRUSTEE is most useful for explaining the decisions of Kitsune NIDS followed by ENIDrift NIDS. Across the datasets, with top-k pruning, Kitsune NIDS yields the highest fidelity values. The comparison of fidelity values obtained during TRUSTEE analysis of the ML-based IoT NIDSs considered in this paper is shown in Table V.
IoT NIDS Fidelity (Without pruning) Fidelity (with top-k pruning, k=10)
Mirai dataset CICIDS-2017 dataset Mirai dataset CICIDS-2017 dataset
Kitsune 0.786 0.843 0.721 0.919
HorusEye -1.365 -0.956 -1.038e-06 -0.018
ENIDrift 0.715 0.929 -1.536e-09 0.53
HELAD 0.0 0.0 0.0 0.0
TABLE V: Comparison of TRUSTEE fidelity values with and without pruning for ML-based IoT NIDSs considered
SHAP evaluation metrics: We experimented with multiple implementations of xAI evaluation metrics (e.g., faithfulness, stability, monotonicity) to compare the kernel SHAP explanations across the four IoT NIDSs under consideration in our paper. Some of those implementations are OpenXAI  [^1] https://github.com/AI4LIFE-GROUP/OpenXAI, XAI-bench  [^2] https://github.com/abacusai/xai-bench and Quantus  [^3] https://github.com/understandable-machine-intelligence-lab/Quantus. However, after extensive experimentation with those implementations, we found out that none of them support custom ML models such as those used for IoT NIDSs. They support out-of-the-box ML models such as linear regression, MLP and ANN only.
VI Discussion and Future Work
From the experimental results obtained in the previous section, we have seen that some ML-based IoT NIDS models (Kitsune, ENIDrift) can be better interpreted in terms of decision trees than other models (HorusEye, HELAD). This might be attributed to the complexity of the NIDS model architecture and the properties of the model components. For example, HorusEye NIDS fidelity values are in the negative. Though HorusEye architecture includes only a single autoencoder in the second stage (which processes control plane traffic) compared to an ensemble of AEs in Kitsune or ENIDrift, that single AE is modified from a conventional AE to keep it lightweight, reduce the computational complexity, reduce the false-positive rate and improve throughput. It uses an asymmetric AE with separable convolution, dilation convolution and model quantization. As it is, AEs are limited in their interpretability as a result of their non-linear nature. Due to the complex architecture of the modified AE deployed in HorusEye, TRUSTEE might not be able to build a DT which can imitate the decisions of the black-box modified AE within an acceptable error rate. Therefore, to build ML-based IoT NIDSs with better “explainability”, we recommend using less complex ML architectures and conventional components. However, there will be a trade-off between “explainability” and model performance (e.g., false-positive rate) and IoT NIDS designers need to achieve a balance between the two.
We have also seen from the results that TRUSTEE and SHAP explanations are in conflict for 2 out of 3 ML-based IoT NIDSs considered in our work. Their explanations are in agreement for HorusEye NIDS but in conflict for Kitsune and ENIDrift NIDSs. This is probably due to the fact that SHAP value for each feature is calculated for a subset of data samples and reflects the importance of that feature in the decision made by the black-box model for a majority of the data samples in that subset. On the other hand, TRUSTEE acts on the complete input dataset and identifies the top features contributing the most to forming the decision boundary separating all benign and malicious data samples which are part of the input dataset. Another reason could be the different optimization criteria used by TRUSTEE and SHAP while building explanations. TRUSTEE applies a teacher-student dynamic derived from imitation learning that uses the black-box ML model as an oracle in conjunction with a carefully curated dataset to guide the training of a surrogate “white-box” model in the form of a DT that imitates the black-box’s decisions. On the other hand, SHAP uses the Shapley kernel method to approximate the black-box model-interpretable model fidelity function which recovers Shapley values. The specific data subset used to conduct the TRUSTEE-SHAP comparison could also play a role in the conflict. To reduce the conflict between TRUSTEE and SHAP explanations at the subset-level, we need to further investigate its underlying cause in the future.
Finally, it can be inferred from the results that some ML-based IoT NIDSs (Kitsune) are more vulnerable to o.o.d. data samples than others (HorusEye, HELAD) while for the rest (ENIDrift), such vulnerability remains inconclusive. One of the possible underlying reasons is that sometimes IoT NIDSs are trained with training data containing higher than real-world attack traffic volume. To reduce vulnerability to o.o.d samples, we recommend that IoT NIDS designers use training data which reflects the attack traffic volume in real-world deployments. Additionally, the NIDS architecture may also have a role to play in its vulnerability to o.o.d. samples. An AE trained on benign traffic may be vulnerable to o.o.d. attack traffic samples as the threshold used for output loss function (RMSE) is determined based on “in-distribution” attack traffic present in the training data. Since, Kitsune consists of an ensemble of AEs, the vulnerability effect is amplified. HorusEye consists of a single modified AE and therefore, its vulnerability to o.o.d. samples is not as pronounced as Kitsune. ENIDrift uses a different method of feature extraction (vector embedding) and updates the weights assigned to AEs in the ensemble based on change in attack traffic distribution which might keep it from producing a discernable pattern of dependence on attack traffic volume compared to Kitsune. We have observed earlier that HELAD NIDS anomaly scores are not vulnerable to o.o.d. samples to a significant extent. This may be attributed to HELAD’s architecture which combines an AE with an LSTM network. Since the final anomaly score is a weighted exponential aggregation of AE’s actual RMSE score and LSTM’s predicted RMSE scores for future samples (using historical samples). Even if the AE is vulnerable to o.o.d. samples, the LSTM’s predicted RMSE scores might be much larger compared to to the AE’s RMSE score (the exponential operation makes the value even larger), thus affecting the final anomaly score to a greater extent and reducing HELAD’s vulnerability to o.o.d. samples.
VII Conclusion
We have analyzed four state-of-the-art ML-based IoT NIDS models (Kitsune, HorsusEye, ENIDRift, HELAD) using explainable AI techniques such as TRUSTEE and SHAP. Using both global explanations and local explanations for the blackbox models’ decisions, we have presented the most prominent features used by each IoT NIDS model considered. We have compared the explanations generated across xAI methods for a given NIDS model (using our devised data subset-level explanation approach) as well as the explanations generated across the NIDS models for a given xAI method (using fidelity and other standard evaluation metrics). Finally, we evaluate the vulnerability of each NIDS model to inductive bias (o.o.d data samples). The results show that: (1) some ML-based IoT NIDS models can be better interpreted in terms of decision trees than other models, (2) TRUSTEE and SHAP explanations are in conflict for most of the IoT NIDS models considered in this work and (3) some IoT NIDS models are more vulnerable to o.o.d samples than other models.
Appendix A More Experimental Results
In this section, we present explanations of the four state-of-the-art ML-based IoT NIDSs (Kitsune, HorusEeye, ENIDrift, HELAD) generated using xAI methods such as TRUSTEE and kernel SHAP where the NIDSs were trained/tested with the CICIDS-2017 dataset.
A-A Case Study- Kitsune $\times$ CICIDS-2017
TRUSTEE Analysis: The results of TRUSTEE analysis for Kitsune are shown in Table VI. Using TRUSTEE with $30\%$ of the original CICIDS-2017 dataset samples and no pruning results in the DT explanation which achieves 0.843 fidelity compared to Kitsune. Using TRUSTEE with $30\%$ of the dataset samples and top-k Pruning method (setting $k=3$) results in a DT explanation that is shown in Fig. 13 and achieves 0.919 fidelity compared to Kitsune.
The top three prominent features that Kitsune uses to determine an anomaly are: $HpHp\_0.1\_pcc\_0\_1$, $HH\_0.1\_pcc\_0\_1$ and $HH\_jit\_0.01\_std$. Thus, Kitsune relies mainly on the sizes of packets and the jitter experienced by packets exchanged per time frame to determine if an attack is underway.
Sample size Top-k pruning used? DT size, depth, leaves Fidelity
30% No 78819, 80, 39410 0.843
30% Yes (k=3) 29, 14, 15 0.919
TABLE VI: TRUSTEE analysis results for Kitsune CICIDS-2017 model
Figure 13: Decision tree for Kitsune CICIDS-2017 model with Top-k pruning (k=3). Only the top 3 layers are shown.
SHAP Analysis: Using kernel SHAP analysis, Fig. 14 shows a summary of how the top features in CICIDS-2017 dataset impact the Kitsune NIDS model’s output through a beeswarm plot for randomly selected 1000 data samples corresponding to the benign root node-leaf node path in the TRUSTEE DT (Fig. 13). Features such as $HpHp\_0.01\_weight\_0$, $HH\_jit\_0.01\_weight$ and $HH\_0.01\_weight\_0$ have the most impact on the prediction value, in that order.
TRUSTEE vs SHAP Analysis: Using the systematic approach for comparing TRUSTEE and SHAP explanations as explained in sub-section IV-B, since the set of features used in the benign root node-leaf node path in the TRUSTEE DT is not a subset of the set of top contributing features computed by kernel SHAP summary analysis ($HpHp\_0.1\_pcc\_0\_1$ is the only common feature between the two sets), TRUSTEE AND SHAP explanations do not agree with each other at the subset level for Kitsune NIDS.
Figure 14: SHAP beeswarm plot for Kitsune CICIDS-2017 model
Detection of Inductive Bias: The CICIDS-2017 dataset does not provide a separate dataset for botnet traffic trace which could be used for comparison with the Mirai trace in terms of the attack traffic volume. However, using a similar approach for tampering as in the case of Mirai trace, we evaluated the following threshold ranges (in packets per second) for ARP requests originating from infected hosts : (1) 10-50, (2) 30-70 and (3) 50-90. The original trace as well as the tampered traces are shown in Fig. 15. It can be observed that: (1) there is no significant difference between the benign and attack traffic volumes for the original trace and (2) tampering had negligible effect on the difference between benign and attack traffic in terms of number of packets per second. Thus, unlike the Mirai trace, there is no significant difference between the benign and attack traffic volumes which can be learnt by the NIDS model and leave it vulnerable to a change in attack traffic distribution. Thus, we do not investigate the presence of inductive bias (vulnerability to o.o.d samples) any further for IoT NIDS models trained with the CICIDS-2017 dataset.
Figure 15: Number of packets per second for original CICIDS-2017 trace and tampered traces. Blue portions of the plot correspond to benign traffic while red portions correspond to malicious traffic inclusive of benign and attack traffic.
A-B Case Study- HorusEye $\times$ CICIDS-2017
TRUSTEE Analysis: The results of TRUSTEE analysis for HorusEye are shown in Table VII. Using TRUSTEE with $30\%$ of the CICIDS-2017 dataset samples with no pruning results in a DT explanation which achieves negative fidelity ($\sim-0.956$) compared to original HorusEye model. Using TRUSTEE with $30\%$ of the dataset samples and top-k Pruning method (setting $k=10$) results in a DT explanation which achieves negative fidelity ($\sim-0.018$) compared to original HorusEye model but higher fidelity compared to no pruning. Negative fidelity values mean that the DT explanation fits HorusEye predictions worse than the mean of the predictions.
Nevertheless, the decision tree for HorusEye Mirai model with top-k pruning (k=10) is shown in Fig. 16. The top three prominent features that HorusEye’s DT uses to determine an anomaly are: $HH\_5\_pcc\_0\_1$, $HpHp\_1\_covariance\_0\_1$ and $HH\_jit\_0.01\_std$. Thus, HorusEye’s DT relies mainly on the sizes of packets and the jitter experienced by packets exchanged between a pair of IP addresses per time frame to determine if an attack is underway.
Sample size Top-k pruning used? DT size, depth, leaves Fidelity
30% No 20265, 94, 10133 -0.956
30% Yes (k=10) 241, 52, 121 -0.018
TABLE VII: TRUSTEE analysis results for HorusEye CICIDS-2017 model
Figure 16: Decision tree for HorusEye CICIDS-2017 model with Top-k pruning (k=10). Only the top 3 layers are shown.
SHAP Analysis: Using kernel SHAP analysis, Fig. 17 shows a summary of how the top features in CICIDS-2017 dataset impact the HorusEye NIDS model’s output using a beeswarm plot for randomly selected 1000 data samples corresponding to the benign root node-leaf node path in the TRUSTEE DT (Fig. 16). Features such as $HH\_jit\_0.1\_std$, $MI\_dir\_0.01\_std$ and $HpHp\_0.1\_weight\_0$ have the most impact on the prediction value, in that order.
Figure 17: SHAP beeswarm plot for HorusEye CICIDS-2017 model
TRUSTEE vs SHAP Analysis: Using the systematic approach for comparing TRUSTEE and SHAP explanations as explained in sub-section IV-B, the set of features used in the benign root node-leaf node path in the TRUSTEE DT is not a subset of the set of top contributing features computed by kernel SHAP summary analysis. Therefore, TRUSTEE AND SHAP explanations do not agree with each other at the subset level for HorusEye NIDS.
A-C Case Study- ENIDrift $\times$ CICIDS-2017
In this analysis, we used the CICIDS-2017-Wed (corresponding to the CICIDS data collected on July 5, 2017 (Wednesday)) dataset for training/testing ENIDrift NIDS.
TRUSTEE Analysis: The results of TRUSTEE analysis for ENIDrift are shown in Table VIII. Using TRUSTEE with $30\%$ of the CICIDS-2017 dataset samples and no pruning results in the DT explanation which achieves 0.929 fidelity compared to ENIDrift. Using TRUSTEE with $30\%$ of the dataset samples and top-k Pruning method (setting $k=10$) results in a DT explanation (shown in Fig. 18) which achieves 0.53 fidelity compared to ENIDrift. This is expected since pruning the branches of a DT reduces its fidelity. The most prominent features ENIDrift uses to determine anomaly are: $f38$, $f81$, $f158$.
Sample size Top-k pruning used? DT size, depth, leaves Fidelity
30% No 253, 24, 127 0.929
30% Yes (k=10) 25, 7, 13 0.53
TABLE VIII: TRUSTEE analysis results for ENIDrift CICIDS-2017 model
Figure 18: Decision tree for ENIDrift CICIDS-2017 model with Top-k pruning (k=10). Only the top 4 layers are shown.
SHAP Analysis: Fig. 19 shows a summary of how the top features in CICIDS-2017 dataset impact the ENIDrift NIDS model’s output using a beeswarm plot for randomly selected 300 data samples. Features (embedding vector components) such as $f31$, $f84$, $f155$ have the most impact on the prediction value, in that order.
Figure 19: SHAP beeswarm plot for ENIDrift CICIDS-2017 model
TRUSTEE vs SHAP Analysis: Using the systematic approach for comparing TRUSTEE and SHAP explanations as explained in sub-section IV-B, the set of features used in the benign root node-leaf node path in the TRUSTEE DT is not a subset of the set of top contributing features computed by kernel SHAP summary analysis (there are no common features between the two sets). Therefore, TRUSTEE AND SHAP explanations do not agree with each other at the subset level for ENIDrift NIDS.
A-D Case Study- HELAD $\times$ CICIDS-2017
In this analysis, we used a subset of the original CICIDS-2017 dataset consisting of 301,000 packets for this analysis as per the open source implementation  [^4] https://github.com/cdogemaru/CPIP released by the authors. The first 200,000 packets were used for training and the rest of the packets were used for testing. Among the packets used for testing, the first 50,000 were benign while the rest were attack-related. The complete CICIDS-2017 dataset was not used since it increased the LSTM training time substantially.
TRUSTEE Analysis: The results of TRUSTEE analysis for HELAD are shown in Table IX. Using TRUSTEE with $30\%$ of the reduced CICIDS-2017 dataset samples and no pruning results in the DT explanation which achieves 0.0 fidelity compared to HELAD. Using TRUSTEE with $30\%$ of the dataset samples and top-k Pruning method (setting $k=10$) results in a DT explanation (shown in Fig. 20) which achieves 0.0 fidelity compared to HELAD.
The top three prominent features HELAD’s DT uses to determine an anomaly are:
  1. 1.
    $HH\_jit\_0.01\_mean$- Packet jitter aggregated by traffic sent between a set of source and destination IP addresses with time window 0.01 (1 minute).
  2. 2.
    $HH\_0.1\_radius\_0\_1$- L2-norm of variances of two packet size streams aggregated by traffic sent between a set of source and destination IP addresses with time window 0.1 (10 seconds).
  3. 3.
    $MI\_dir\_1.0\_std$- Standard deviation of packet sizes aggregated by source MAC and IP addresses, with time window 1.0 (1.5 seconds).
Thus, HELAD’s DT relies mainly on the jitter experienced and the sizes of packets exchanged between a pair of IP addresses per time frame to determine if an attack is underway.
Sample size Top-k pruning used? DT size, depth, leaves Fidelity
30% No 19607, 94, 9804 0.0
30% Yes (k=10) 475, 77, 238 0.0
TABLE IX: TRUSTEE analysis results for HELAD NIDS
Figure 20: Decision tree for HELAD CICIDS-2017 model with Top-k pruning (k=10). Only the top 3 layers are shown.
SHAP Analysis: Running a kernel SHAP analysis of HELAD NIDS with the reduced CICIDS-2017 dataset resulted in issues. First, we ran the analysis with 200,000 packets but it threw a numpy memory error since it was unable to allocate a huge amount of memory (313 GiB) for storing a ($2248\times 20000000$) array with 64-bit float values. Next, we ran the analysis with 20,000 packets but the LSTM was taking too long to train. Hence, we decided to not proceed with the kernel SHAP analysis.